for a dummy variable regression, female = 0, male =1......if we accidentally code the female = -1 instead of 0, what's the estimate of OLS of beta0 and beta1?
Answer #1

Solution: For a dummy variable regression, female = 0, male =1,if we accidentally code the female = -1 instead of 0, the estimate of OLS of \beta_{0} and \beta_{1} will be

Lets denote the estimate of OLS of beta0 and beta1 with correct coding that is female = 0, male =1

beta0 denoted by b_{0}

and beta1 denoted by b_{1}.

And the estimate of OLS of beta0 and beta1 with incorrect coding that is female = -1, male =1

beta0 denoted by \beta_{0}

and beta1 denoted by \beta_{1}.

Therefore the esimate of OLS of \beta_{0} and \beta_{1} will be:

\beta_{0}=b_{0}+\left ( \frac{b_{1}}{2} \right )

\beta_{1}=\left ( \frac{b_{1}}{2} \right )